# Retraining-Free Merging of Sparse Mixture-of-Experts via Hierarchical Clustering

### v1

@@ 3.2.3 Expert Merging
Upon completion of clustering, HC-SMoE proceeds to merge the experts within each cluster, as illustrated in Fig. 2 (c). Our empirical evidence indicates that while the choice of merging method does influence the overall performance, its impact is relatively modest compared to the significance of robust clustering results. This study considers three merging strategies: average merging, frequency-weighted merging, and fixed-dominant merging. Average merging computes the element-wise average of the weights of all experts in each cluster. Frequency-weighted merging (Li et al., 2024) is performed as $E_{merged}=\frac{\sum_{i=1}^{k}\alpha_{i}E_{i}}{\sum_{i=1}^{k}\alpha_{i}}$ for a group of $k$ experts $E_{1},...,E_{k}$, where $\alpha_{i}$ denotes the usage frequency of expert $E_{i}$. Fixed-dominant merging, a method introduced in this study, represents an efficient adaptation of ZipIt specifically designed for merging experts in SMoE models.
In fixed-dominant merging, the expert closest to the cluster center is selected as the dominant expert, which serves as a reference point. The features of the other experts in the cluster are then re-ordered based on their correlation with the dominant expert’s features. Subsequently, an averaging operation is performed between the dominant expert and the re-ordered features of the remaining experts. This method is specifically tailored for SMoEs, as it preserves the structural characteristics of the dominant expert while allowing efficient merging. Unlike the original ZipIt algorithm, which requires exhaustive pairwise comparisons, our adaptation is computationally efficient and ensures the preservation of key characteristics of the dominant expert, making it highly suitable for the SMoE framework. An ablation study for different merging approaches is presented in Section 4.3. For elaboration on fixed-dominant merging and comparison to ZipIt, please refer to Appendix A.2.
@@ 4 Experimental Results
In the following sections, we first describe the experimental setups in Section 4.1. Next, we compare the performance of HC-SMoE against various SMoE model pruning and merging baselines. Finally, we present an ablation study to assess the effectiveness of each component in HC-SMoE’s design.
4.1 Experimental Settings
To evaluate our method in a task-agnostic setting, we utilize eight tasks using the EleutherAI Language Model Evaluation Harness (Gao et al., 2024). These are designed to cover various aspects of language understanding and reasoning, including both Challenge and Easy sets in AI2 Reasoning Challenge (ARC) (Clark et al., 2018), BoolQ (Clark et al., 2019), HellaSwag (Zellers et al., 2019), Massive Multitask Language Understanding (MMLU) (Hendrycks et al., 2021), OpenBookQA (Mihaylov et al., 2018), Recognizing Textual Entailment (RTE) (Bentivogli et al., 2009) and Winograd Schema Challenge (Sakaguchi et al., 2021). We report zero-shot accuracy on those benchmarks.
For our comparisons, three pruning baselines are employed: O-prune (Lu et al., 2024), S-prune (He et al., 2024), and F-prune. F-prune, where ‘F’ denotes frequency, adheres to the same methodology as S-prune. However, it employs frequency as the criterion for pruning experts, in contrast to S-prune which utilizes router logits. Due to the high computational complexity of O-prune on Qwen, a random sampling of $10,000$ possible expert sets in each layer is performed instead. The set with the smallest output difference from the original model is selected, denoted as O-prune ($10^{5}$) in the Qwen experiments. In addition, M-SMoE is included as the merging baseline and applied in a task-agnostic setting without retraining to ensure a fair comparison. All baselines and HC-SMoE require a calibration dataset to estimate input statistics. This dataset is constructed by sampling from the C4 corpus (Raffel et al., 2020), concatenating extracted text into $32$ sequences of $2,048$ tokens each.
4.2 Performance Comparisons
Table 2: Zero-shot performance evaluation of different expert pruning and merging methods on Qwen with reducing experts to 45 and 30 per layer. HC-SMoE (avg) stands for average linkage method when performing hierarchical clustering. HC-SMoE (single) stands for single linkage.
@@ 4.2 Performance Comparisons
Table 3: Zero-shot performance evaluation of different expert pruning and merging methods on Mixtral 8x7B with reducing experts to six and four per layer.
@@
This section presents a comprehensive comparison of the performance of the models reduced by the proposed HC-SMoE against the original SMoE models and the baselines. The analysis encompasses various model sizes and tasks, and provides insights into the efficacy and scalability of the proposed HC-SmoE method. As presented in Tables 2 and 3, the M-SMoE baseline exhibits the lowest performance across all benchmarks, indicating the ineffectiveness of router-logit-based grouping in a task-agnostic setting. O-prune demonstrates suboptimal performance, particularly on Qwen, due to its limitations in evaluating all possible expert combinations. This results in a substantial performance decline compared to Mixtral. In contrast, HC-SMoE demonstrates consistent superiority over these baselines, irrespective of model size, and proves applicable to different number of experts.
It is noteworthy that Qwen 45x2.7B and Mixtral 4x7B achieve comparable scores despite a twofold difference in parameter count. This observation substantiates the scalability of HC-SMoE to SMoE models with a higher number of experts. With a 25% reduction in experts, our method even surpasses the original model on certain tasks, such as Mixtral 6x7B on BoolQ and Qwen 45x2.7B on RTE. This improvement can be attributed to the reduction of expert redundancy after merging. In this configuration, both Qwen and Mixtral exhibit an average performance gap of less than 3% compared to their original models. Even with a 50% reduction, HC-SMoE applied to Qwen maintains a gap of merely 7.43% and outperforms the best baseline, F-prune, which lags behind HC-SMoE by 7.46%. These results validate the robustness and efficacy of HC-SMoE across diverse model sizes and tasks.
@@ 4.3 Ablation Study
Table 4: Different linkage method comparisons of hierarchical clustering on Qwen 45x2.7B.
@@
expert-output 0.3592 0.7942 0.270 0.7292 0.5382
Complete router-logits 0.3677 0.7694 0.248 0.7329 0.5295
@@
expert-output 0.2338 0.6037 0.210 0.5162 0.3909
Average router-logits 0.2073 0.3801 0.172 0.5018 0.3153
@@
expert-output 0.3660 0.7948 0.272 0.7509 0.5459
Table 5: Our HC-SMoE compares to K-means clustering. K-means-fix assigns the first $r$ experts as initial center. K-means-rnd randomly choose $r$ experts as initial center. For each task, we highlight the best performance in yellow, and mark the best performance within same cluster method bold.
Model Cluster Metric ARC-c BoolQ OBQA RTE Average
Qwen 60x2.7B None None 0.3951 0.8135 0.310 0.7329 0.5629
Qwen 45x2.7B K-means-fix router-logits 0.3669 0.7664 0.270 0.6679 0.5178
weight 0.3797 0.7294 0.268 0.6751 0.5131
expert-output 0.3925 0.7850 0.270 0.7184 0.5415
K-means-rnd router-logits 0.3797 0.7621 0.276 0.6029 0.5052
weight 0.2432 0.5557 0.154 0.5812 0.3835
expert-output 0.3797 0.7177 0.270 0.6968 0.5161
HC-SMoE expert-output 0.3646 0.7927 0.268 0.7449 0.5426
Qwen 30x2.7B K-means-fix router-logits 0.2031 0.4015 0.162 0.4838 0.3126
@@
expert-output 0.3370 0.6398 0.224 0.6065 0.4518
HC-SMoE expert-output 0.3515 0.7544 0.228 0.6631 0.4993
Table 6: Comparisons for different similarity metric to single-shot grouping method and our HC-SMoE on Mixtral 8x7B with reducing experts to average 6 and 4 per layer.
@@
weight 0.4974 0.7955 0.781 0.6131 0.5244 0.34 0.6715 0.7585 0.6227
expert-output 0.506 0.8056 0.8373 0.613 0.5595 0.306 0.6318 0.7474 0.6258
@@
weight 0.4172 0.7382 0.7862 0.5457 0.4223 0.256 0.5523 0.7143 0.554
expert-output 0.4326 0.7386 0.8021 0.5467 0.429 0.278 0.5704 0.7245 0.5652
@@
Single-shot Grouping v.s. Hierarchical Clustering. In this analysis, we follow the single-shot grouping methods outlined in Li et al. (2024) to compare results on Mixtral 8x7B, and report the results in Table 6. Among the similarity metrics evaluated, router-logits exhibits the poorest performance, indicating its unsuitability for task-agnostic settings due to its reliance on dataset-specific statistics. In both the 25% and 50% parameter reduction scenarios, all one-shot grouping methods underperform compared to O-prune presented in Table 3. This observation suggests that these grouping methods fail to form effective clusters, and can potentially result in lower performance even when attempting to absorb all expert knowledge. The method based on the expert output metric demonstrates superior performance over other similarity metrics. It outperforms router-logits by 24.01% and weights by 1.12% when reducing 50% of the expert parameters. This finding highlights the importance of selecting appropriate similarity metrics for effective expert grouping. HC-SMoE demonstrates a clear advantage over the one-shot grouping approaches. It achieves average improvements of 1.98% and 1.67% in the 25% and 50% parameter reduction settings, respectively.
Table 7: Various merging methods with hierarchical clustering average linkage based on expert outputs. Fix-Dom represents fixed-dominant merging described in Section 3.2.3. Avg in the Merge column denotes the average score among all the merging strategy under same model settings.
@@ 5 Conclusion
Appendix A Exploratory Experiments
All the experiments code is released at https://github.com/wazenmai/HC-SMoE.
A.1 Non-Uniform Hierarchical Clustering
@@
In our main experiments, the number of clusters in each layer is fixed and uniform due to model design choices. Here, we explore a more flexible approach that allows different numbers of clusters in each layer while maintaining an overall 25% or 50% reduction of experts. To determine the cluster count per layer, we first select the top $r$most frequently activated experts based on their activation frequencies across layers. We then count the number of these experts that remain in each layer to guide the selection of clusters in that layer, followed by hierarchical clustering.
@@
For example, in the uniform clustering setting for Qwen with a 25% reduction, the distribution will be $[45,45,45,45,...,45]$ across all layers. In contrast, the non-uniform setting might result in a distribution like $[48,45,40,42,50,...]$, as long as the overall number of clusters aligns with the target reduction. Table 8 presents the results of this non-uniform clustering strategy.
Table 8: The comparison between ZipIt and Fix-Dom merging for reducing 25% experts of Qwen under the same expert clustering groups.
@@
Fix-Dom 0.3558 0.6582 0.7917 0.5558 0.5306 0.270 0.7256 0.6559 0.5680
A.2 Fixed-Dominant Merging

### v4

@@ 3.2.3 Expert Merging
Following cluster formation, HC-SMoE merges experts within each cluster via a parametrized weight-space aggregation. Our empirical evidence indicates that while the choice of merging method does influence the overall performance, its impact is relatively modest compared to the significance of clustering results. Specifically, HC-SMoE merges the clustered experts on the weight space as: $\hat{E}_{i}=\sum_{j=1}^{|C_{i}|}\alpha_{j}E_{j},\text{where}\quad\sum_{j=1}^{|C_{i}|}\alpha_{j}=1$, and $\alpha_{j}$ denotes the weight for merging expert $j$. This study considers three different merging strategies: average merging, frequency-weighted merging, as well as fixed-dominant merging. In average merging, $\alpha_{j}=\frac{1}{|C_{i}|}$. In frequency-weighted merging, $\alpha_{j}$ denotes the usage frequency of expert $j$. On the other hand, fixed-dominant merging, a methodology introduced in this study, represents an efficient adaptation of ZipIt specifically developed for merging experts in SMoE models. As developing novel merging methods extends beyond the focus of this article, we present the comprehensive analysis of fixed-dominant merging and comparisons with ZipIt in Appendix B.2. HC-SMoE implements frequency-weighted merging to maintain merging strategy flexibility, while preserving the task-agnostic nature of hierarchical clustering. The experimental results presented in Table 7 further reveal that merging strategy selection has marginal impact when utilizing a general-purpose calibration dataset.
@@ 4 Experimental Results
Table 2: Zero-shot comparison of Qwen1.5-MoE-A2.7B-Chat: original architecture v.s. reduced versions with 45 and 30 experts per layer. HC-SMoE (avg) stands for average linkage when performing hierarchical clustering. HC-SMoE (single) stands for single linkage.
@@
4.1 Experimental Settings
To evaluate our method in a task-agnostic setting, we utilize eight tasks using the EleutherAI Language Model Evaluation Harness (Gao et al., 2024). These are designed to cover various aspects of language understanding and reasoning, including both Challenge and Easy sets in AI2 Reasoning Challenge (ARC) (Clark et al., 2018), BoolQ (Clark et al., 2019), HellaSwag (Zellers et al., 2019), Massive Multitask Language Understanding (MMLU) (Hendrycks et al., 2021a), OpenBookQA (Mihaylov et al., 2018), Recognizing Textual Entailment (RTE) (Bentivogli et al., 2009) and Winograd Schema Challenge (Sakaguchi et al., 2021). We report zero-shot accuracy on those benchmarks. To demonstrate HC-SMoE’s domain adaptability, additional evaluations on medical reasoning tasks (are offered in Appendix B.4.2).
For our comparisons, three pruning baselines are employed: O-prune (Lu et al., 2024), S-prune (He et al., 2024), and F-prune. F-prune, where ‘F’ denotes frequency, adheres to the same methodology as S-prune. However, it employs frequency as the criterion for pruning experts, in contrast to S-prune which utilizes router logits. Due to the high computational complexity of O-prune on Qwen, a random sampling of $10,000$ possible expert sets in each layer is performed instead. The set with the smallest output difference from the original model is selected, denoted as O-prune ($10^{5}$) in the Qwen experiments. In addition, M-SMoE is included as the merging baseline and applied in a task-agnostic setting without retraining to ensure a fair comparison. All baselines and HC-SMoE require a calibration dataset to estimate input statistics. This dataset is constructed by sampling from the C4 corpus (Raffel et al., 2020), concatenating extracted text into $32$ sequences of $2,048$ tokens each. To further validate the independence of HC-SMoE from the calibration dataset, we construct two additional datasets from MATH (Hendrycks et al., 2021b) and CodeQA (Liu & Wan, 2021). Please refer to our Appendix B.3 for more details.
4.2 Performance Comparisons
Table 3: Zero-shot comparison of Mixtral 8x7B: original architecture v.s. reduced versions with six and four experts per layer.
@@ 4.2 Performance Comparisons
This section presents a comprehensive comparison of the performance of the models reduced by HC-SMoE against the original SMoE models and the baselines. The analysis encompasses various model sizes and tasks, and provides insights into the efficacy and scalability of the proposed HC-SmoE method. As presented in Tables 2 and 3, the M-SMoE baseline exhibits the lowest performance across all benchmarks, indicating the ineffectiveness of router-logit-based grouping in a task-agnostic setting. O-prune demonstrates suboptimal performance, particularly on Qwen, due to its limitations in evaluating all possible expert combinations. This results in a substantial performance decline compared to Mixtral. In contrast, HC-SMoE demonstrates consistent superiority over these baselines, irrespective of model size, and proves applicable to different numbers of SmoE experts.
Qwen 45x2.7B and Mixtral 4x7B achieve comparable scores despite a twofold difference in parameter count. This observation substantiates the scalability of HC-SMoE to SMoE models with a higher number of experts. With a 25% reduction in experts, our method even surpasses the original model on certain tasks, such as Mixtral 6x7B on BoolQ and Qwen 45x2.7B on RTE. This improvement can be attributed to the reduction of expert redundancy after merging. In this configuration, both Qwen and Mixtral exhibit an average performance gap of less than 3% compared to their original models. Even with a 50% reduction, HC-SMoE applied to Qwen maintains a gap of merely 7.43% and outperforms the best baseline, F-prune, which lags behind HC-SMoE by 7.46%. The experimental results validate the robustness and efficacy of HC-SMoE over diverse model sizes and tasks.
@@ 4.3 Ablation Study
Table 4: Different linkage method comparisons of hierarchical clustering on Qwen 45x2.7B, where ‘rl’ denotes router-logits and ‘eo’ denotes expert-output.
@@
eo 0.3592 0.7942 0.270 0.7292 0.5382
Complete rl 0.3677 0.7694 0.248 0.7329 0.5295
@@
eo 0.2338 0.6037 0.210 0.5162 0.3909
Average rl 0.2073 0.3801 0.172 0.5018 0.3153
@@
eo 0.3660 0.7948 0.272 0.7509 0.5459
Table 5: Performance comparison between the proposed HC-SMoE and K-means clustering approaches on Qwen 30x2.7B. K-means-fix designates the first $r$ experts as initial centers, while K-means-rnd randomly selects $r$ experts as initial centers. Note that the best performance across all methods are marked in bold, with the best performance in each clustering method marked with underline.
Cluter Metric ARC-c BoolQ OBQA RTE Average
K-fix rl 0.2031 0.4015 0.162 0.4838 0.3126
@@
eo 0.3370 0.6398 0.224 0.6065 0.4518
HC eo 0.3532 0.7535 0.228 0.6606 0.4988
Table 6: Comparisons for different similarity metric to single-shot grouping method and our HC-SMoE on Mixtral 8x7B when reducing SMoE experts to average six and four per layer.
@@
weight 0.4974 0.7955 0.7810 0.6131 0.5244 0.340 0.6715 0.7585 0.6227
expert-output 0.5060 0.8056 0.8373 0.6130 0.5595 0.306 0.6318 0.7474 0.6258
@@
weight 0.4172 0.7382 0.7862 0.5457 0.4223 0.256 0.5523 0.7143 0.5540
expert-output 0.4326 0.7386 0.8021 0.5467 0.4290 0.278 0.5704 0.7245 0.5652
@@
Table 7: Various merging methods with HC average linkage based on expert outputs. Fix-Dom represents fixed-dominant merging described in Section 3.2.3. Avg in the Merge column denotes the average score among all the merging strategy under same model settings.
@@
Single-shot Grouping v.s. Hierarchical Clustering. In this analysis, we follow the single-shot grouping methods outlined in (Li et al., 2024) to compare results on Mixtral 8x7B, and report the results in Table 6. Among the similarity metrics evaluated, router-logits exhibits the poorest performance, indicating its unsuitability for task-agnostic settings due to its reliance on dataset-specific statistics. In both the 25% and 50% parameter reduction scenarios, all one-shot grouping methods underperform compared to O-prune presented in Table 3. This observation suggests that these grouping methods fail to form effective clusters, and can potentially result in lower performance even when attempting to absorb all expert knowledge. The method based on the expert output metric demonstrates superior performance over other similarity metrics. It outperforms router-logits by 24.01% and weights by 1.12% when reducing 50% of the expert parameters. This finding highlights the importance of selecting appropriate similarity metrics for effective expert grouping. The results reveal that HC-SMoE demonstrates a clear advantage over the one-shot grouping approaches. It achieves average improvements of 1.98% and 1.67% in the 25% and 50% parameter reduction settings, respectively.
@@ 5 Conclusion
Impact Statement
In compliance with ICML’s guidelines, we affirm that our HC-SMoE algorithm does not involve human subjects, personal data, or any components that could raise ethical concerns. The algorithm focuses solely on clustering machine learning model experts based on output similarity to optimize parameter efficiency, ensuring no ethical issues are present.
Acknowledgements
The authors gratefully acknowledge support from the National Science and Technology Council (NSTC) in Taiwan under grant numbers MOST 111-2223-E-002-011-MY3, NSTC 113-2221-E-002-212-MY3, and NSTC 113-2640-E-002-003. We also express our sincere appreciation to NVIDIA Corporation and the NVIDIA AI Technology Center (NVAITC) for the donation of GPUs and access to the Taipei-1 supercomputer. Furthermore, we thank the National Center for High-Performance Computing for providing computational and storage resources.
Appendix: Retraining-Free Merging of Sparse MoE via Hierarchical Clustering
Appendix A Theoretical Justification
In this section, we present a theoretical rationale for the HC-SMoE algorithm, which employs hierarchical clustering with unweighted average linkage to group experts into clusters and reduce the number of parameters in a Mixture-of-Experts (MoE) model. Our objective is to minimize the approximation error between the original MoE output and the HC-SMoE output while leveraging the performance guarantees for hierarchical clustering.
Let $\{G_{j}\}_{j=1}^{r}$ denote a partition sets of the expert indices into $r$ clusters. We define the average-merged expert for group $j$ as:
$\bar{E}_{j}(x):=\frac{1}{|G_{j}|}\sum_{i\in G_{j}}E_{i}(x),$ (9)
where $i$ represents the expert index and $g(i)$ is the function that maps expert $E_{i}$ to its cluster index $j=g(i)$. The HC-SMoE output can be expressed as:
$y_{\mathrm{HC}}(x)=\sum_{j=1}^{r}(\sum_{i\in G_{j}}P_{i}(x))\cdot\bar{E}_{j}(x)=\sum_{i=1}^{n}P_{i}(x)\cdot\bar{E}_{g(i)}(x),$ (10)
where $P_{i}(x)$ denotes the routing probability assigned to expert $E_{i}$ for input $x$, and $n$ represents the total number of experts. The approximation error between the original MoE output and the HC-SMoE output is formulated as:
$\|y_{\mathrm{orig}}(x)-y_{\mathrm{HC}}(x)||^{2}=||\sum_{i=1}^{n}P_{i}(x)(E_{i}(x)-\bar{E}_{g(i)}(x))||^{2}\leq\sum_{i=1}^{n}P_{i}(x)\cdot||E_{i}(x)-\bar{E}_{g(i)}(x)\|^{2},$ (11)
by Jensen’s inequality applied to the convex function $\|\cdot\|^{2}$. This inequality demonstrates that the HC-SMoE approximation error is bounded by the weighted intra-cluster variance. To reduce the approximation error $\|y_{\mathrm{orig}}(x)-y_{\mathrm{HC}}(x)\|^{2}$, we opt to to minimize the intra-cluster variance $\|E_{i}(x)-\bar{E}_{g(i)}(x)\|^{2}$.
A.1 Clustering Optimality and Approximation Bound
Finding the optimal solution, defined as $\mathrm{OPT}:=\min\|E-\bar{E}\|^{2}$, is an NP-hard problem. A naive approach would involve exhaustively evaluating all possible partitions of the $n$ experts into $r$ clusters, where the total number of such partitions is given by:
$\frac{1}{r!}\sum_{i=0}^{r}(-1)^{i}\binom{r}{i}(r-i)^{n}.$ (12)
Hierarchical clustering with average linkage is a polynomial-time algorithm that provides an approximation to the optimal clustering. In the worst case, it guarantees a solution satisfying $\|E-\bar{E}\|^{2}\leq 3\cdot\mathrm{OPT}$ (Moseley & Wang, 2023). Therefore, we adopt this computationally efficient method in place of the naive exhaustive approach.
Furthermore, unlike K-means and K-center which depend on random initialization, HC operates as a deterministic algorithm. This deterministic nature ensures stable results across executions. Our empirical evidence indicates that HC with average linkage performs effectively in practice, and is able to deliver robust clustering results without randomness requirements, which benefits parameter reduction in MoE models.
A.2 HC-SMoE Algorithm
The HC-SMoE algorithm performs hierarchical clustering on expert outputs and merges experts within each cluster based on frequency-weighted averaging.
Algorithm 1 HC-SMoE: Hierarchical Clustering for Sparse Mixture-of-Experts
0: Calibration dataset $\mathcal{D}_{cal}$ with $T$ tokens, number of experts within an MoE layer $n$, set of experts $\mathbb{E}=\{E_{1},\dots,E_{n}\}$, target number of clusters $r$.
0: Merged expert set $\{\bar{E}_{1},...,\bar{E}_{r}\}$.
1:// Collect averaged expert outputs
2:for each expert $E_{i}\in\mathbb{E}$ do
3:  $o_{i}=\mathbb{E}_{x\sim\mathcal{D}_{cal}}[E_{i}(x)]$ // Compute mean expert output
4:end for
5:// Hierarchical clustering
6: Initialize clusters $\mathcal{C}=\{\{E_{1}\},\{E_{2}\},...,\{E_{n}\}\}$ and set $m\leftarrow n$
7:while $m>r$ do
8:  Compute pairwise Euclidean distance between expert outputs:
$d(C_{a},C_{b})=\frac{1}{|C_{a}||C_{b}|}\sum_{o_{i}\in C_{a},o_{j}\in C_{b}}\|o_{i}-o_{j}\|_{2}$
9:  Merge two clusters $(C_{a},C_{b})$ with smallest $d(C_{a},C_{b})$ using average linkage.
10:  Update $\mathcal{C}\leftarrow\mathcal{C}\setminus\{C_{a},C_{b}\}\cup\{C_{a}\cup C_{b}\}$ and $m\leftarrow m-1$.
11:end while
12:// Frequency-weighted merging
13:for each cluster $C_{j}\in\mathcal{C}$ do
14:  Let $f_{i}$ be the frequency of tokens assigned to expert $E_{i}$.
15:  Normalize: $\tilde{f}_{i}=\frac{f_{i}}{\sum_{E_{i}\in C_{j}}f_{i}}$.
16:  Compute merged expert:
$\bar{E}_{j}=\sum_{E_{i}\in C_{j}}\tilde{f}_{i}E_{i}$
17:end for
Appendix B Exploratory Experiments
B.1 Non-Uniform Hierarchical Clustering
@@
In our main experiments, the number of clusters in each layer is fixed and uniform due to model design choices. Here, we explore a more flexible approach that allows different numbers of clusters in each layer while maintaining an overall 25% or 50% reduction of experts. To determine the cluster count per layer, we first select the top rr% most frequently activated experts based on their activation frequencies across layers. We then count the number of these experts that remain in each layer to guide the selection of clusters in that layer, followed by hierarchical clustering.
@@
For example, in the uniform clustering setting for Qwen with a 25% reduction, the distribution will be [45,45,45,45,,45][45,45,45,45,...,45] across all layers. In contrast, the non-uniform setting might result in a distribution like [48,45,40,42,50,][48,45,40,42,50,...], as long as the overall number of clusters aligns with the target reduction. Table 8 presents the results of this non-uniform clustering strategy.
Table 8: Zero-shot performance evaluation of non-uniform hierarchical clustering for reducing 25% experts of Qwen. We present the clustering results under single and average linkage with weight and expert-output as similarity metric.
@@
Fix-Dom
0.3558 0.6582 0.7917 0.5558 0.5306 0.270 0.7256 0.6559 0.5680
B.2 Fixed-Dominant Merging